OpenStudy (anonymous):

help solving inequalities? x – 7 > –3.1

OpenStudy (anonymous):

x-7>-3.1 x-7+7>-3.1+7 x>3.9

OpenStudy (anonymous):

x>39/10

OpenStudy (anonymous):

39/10 = 3.9 It is the same

OpenStudy (anonymous):

3.9 =3.9/1 =((3.9)(10))/((1)(10)) <--- multiply numerator and denominator by 10 at the same time =39/10